Circular No. 8270

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 WT_{42} (LINEAR) The Central Bureau has received word that a weak coma has been

imaged for 2003 WT_{42}, an object originally reported as asteroidal by LIN-EAR (discovery observation below; cf. MPEC 2003-W48, MPS 92017), by R. P. Binzel (at the Kitt Peak 4-m telescope) and by J. Licandro, M. Serra-Ricart, J. de Leon Cruz, and N. Pinilla-Alonso (at the 3.56-m Telescopio Nazionale Galileo + Near Infrared Camera-Spectrograph and the 2.5-m Nordic Optical Telescope + Andalucia Faint Object Spectrograph and Camera). Binzel reports that, in 1′′.5 seeing on 2003 Dec. 29.1–29.2 UT with the TV guider, 2003 WT_{42} appeared distinctly more diffuse than stars of similar brightness, with a coma diameter of ∼ 2′′; broadband (500– 900 nm) images showed larger north-south FWHM profiles when compared to stars of similar brightness. Licandro et al. report that a coma diameter of ∼ 6′′–10′′ (total mag R = 17.4 ± 0.1; R − J = 0.8 ± 0.15, which is close to the solar color) was clearly seen on simultaneous infrared and visible (broadband R and J_{s}) images of 2003 WT_{42} obtained on 2004 Jan. 14.9.

2003 UT  α_{2000}  δ_{2000}  Mag.
Nov. 19.26014  4^{h}04^{m}45.^{s}90  +31^{o}43′23′′.0  17.5

Additional astrometry, the following orbital elements, and an ephemeris appear on MPEC 2004-B02.

Epoch = 2006 Apr. 15.0 TT

T = 2006 Apr. 11.1570 TT  ω = 92.^{o}4652
e = 1.001833  Ω = 48.4575 }2000.0
q = 5.192305 AU  i = 31.4164

SUPERNOVA 2004C IN NGC 3683 W. P. S. Meikle, Imperial College, London; and S. Mattila, Stockholm

Observatory, report observations of SN 2004C carried out by D. Carter and R. J. Smith, Liverpool John Moores University, on Jan. 16.2 UT using the Liverpool Telescope (+ RATCam) on La Palma. Preliminary photometry yields the following values: g′ = 18.31, r′ = 16.74, i′ = 16.37, z′ = 16.12 (uncertainties ∼ ± 0.05 mag). Comparison of the g′ − r′ and r′ − i′ colors with the intrinsic g − r and r − i values for a type-Ic supernova (Poznanski et al. 2002, PASP 114, 833, Fig. 8) at 3 weeks past maximum light (cf. IAUC 8269) indicates an extinction exceeding A_{v} = 1.0.

2004 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green